OpenStudy (anonymous):

Can someone help me with this math problem. Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2) I'm having a hard time how to input those into the formula.

OpenStudy (anonymous):

midpoint of x= (x2 + x1)/2 = (-2 + -3)/2 midpoint of y= (y2 + y1)/2 = (1 + 2)/2

OpenStudy (rulnick):

The x value of the midpoint has the average of the two x values. Same for the y's.

OpenStudy (anonymous):

where were you having a hard time?

OpenStudy (anonymous):

@ abb50 ; ah i see where you're putting those. i was mixing them up .. thank you!

OpenStudy (anonymous):

midpoint = x2 + x1 whole thing divided by 2, and y2 + y1 whole thing divided by 2, giving you the midpoint.